# Pinning control of chimera states in systems with higher-order interactions

### v1

@@ II.1 Stuart-Landau oscillators coupled via nonlocal hyperrings
$\begin{cases}\displaystyle\dot{x}_{i}=\alpha x_{i}-\omega y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)x_{i}+\varepsilon\sum_{j}^{N}A_{i,j}\Big{(}x_{j}^{3}-x_{i}^{3}\Big{)},\\ \\ \displaystyle\dot{y}_{i}=\omega x_{i}+\alpha y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)y_{i}+\varepsilon\sum_{j}^{N}A_{i,j}\Big{(}x_{j}^{3}-x_{i}^{3}\Big{)}.\end{cases}$ (7)
The equations for interactions of different orders can be constructed analogously by means of the coupling functions (5) (see also the SM of Muolo et al. (2024)).
@@ II.2 Hyperedge-based local order parameter
To quantify the synchronization of an ensemble of oscillators it is common to use the Kuramoto order parameter Kuramoto (1975), which gives a global measure of how much the oscillators are synchronized. However, chimera states involve the coexistence of coherent (i.e., synchronized) and incoherent (i.e., desynchronized) oscillators and the respective regions are localized. Hence, a global measure of the synchronization does not provide useful information on the chimera state. For this reason, scholars have proposed a local Kuramoto order parameter, which measures the synchronization in a given region of the network, by quantifying the differences between neighboring oscillators, as was done, for instance, in Gambuzza and Frasca (2016). In the case of hyperrings, a natural extension of the local order parameter would need to account for the synchronization inside each hyperedge rather than some arbitrary neighborhood. Partially inspired by a work by Shanahan Shanahan (2010), where the order parameter is defined with respect to the communities of a network, we hereby define a hyperedge-based local order parameter as follows:
@@
where $\mathrm{\imath}$ is the imaginary unit, $\vartheta_{j}$ is the (time evolving) phase of the $j$-th oscillator and $\mathcal{E}_{i}$ is the hyperedge(s) node $i$ is part of. By taking as example a $3$-hyperring, Fig. 1a), we can see that, if node $i$ is a junction node, then it is part of $2$ different hyperedges and will have $6$ neighboring nodes, while non-junction nodes will have only $3$ neighbors. From that, we can observe that non-junction nodes will have the same hyperedge-based local order parameter $\mathcal{R}_{i}^{\mathcal{E}}$. The number of nodes with the same $\mathcal{R}_{i}^{\mathcal{E}}$ in each hyperedge increases with the order of the hyperring: for instance, in $3$-hyperrings they will be $2$, while in $6$-hyperrings they will be $5$.
@@ II.3 Pinning control
Chimera states are often elusive patterns, emerging only for limited ranges of the parameters and specific initial conditions Zakharova (2020). Higher-order interactions greatly enhance the possibility of observing such a behavior Kundu and Ghosh (2022b); Li et al. (2023); Muolo et al. (2024). Nonetheless, ad hoc initial conditions remain a fundamental prerequisite for the chimera to emerge. Our goal is, hence, to induce chimera states in settings where they would not spontaneously appear. For this, we put to use a popular technique in control theory, called pinning control, which consists in externally acting on a subset of the nodes to drive the dynamics of the whole ensemble of nodes towards a desired state Grigoriev et al. (1997); Sorrentino et al. (2007), and has been successfully applied in the context of chimera states on networks Gambuzza and Frasca (2016). In our setting, the pinning will act as a perturbation on a subset of the nodes, with the goal of developing a region of incoherence, while leaving the unperturbed nodes in their synchronous state.
Figure 2: Scheme of the pinning control for a $3$-hyperring of $24$ nodes. The system starts in the synchronized state and the control input is applied to a subset of the nodes to induce the emergence of a chimera state.
In order to implement the control, we need to determine which nodes to pin and for how much time. Hence, let us define the pinning time $t_{p}$, i.e., the time during which the control will be active, and the number of pinned nodes, $N_{p}<N$. We will denote all the nodes with the index $i=1,...,N$ and the subset of pinned nodes with $i_{p}=1,...,N_{p}$. Let us observe that, in the context of pairwise interactions, a chimera state is obtained when about half of the nodes are controlled Gambuzza and Frasca (2016). In Sec. III, where we show the numerical results, the reader will appreciate the great advantage offered by the presence of higher-order interactions. The pinning control setting is schematically depicted in Fig. 2.
@@ II.3.1 Control protocol I: additive pinning control
Let us now proceed with setting up the control protocol. The first type of pinning control is called additive, and it was successfully implemented in Gambuzza and Frasca (2016) to control chimera states with pairwise interactions. In such a setting, the pinned nodes receive an input in an additive fashion. The equations read
$\dot{\vec{X}}_{i}=\vec{F}(\vec{X}_{i})+\mathbb{D}\sum_{j_{1},...,j_{d}=1}^{N}A_{i,j_{1},...,j_{d}}^{(d)}\Big{(}\vec{H}(\vec{X}_{j_{1}},...,\vec{X}_{j_{d}})-\vec{H}^{(d)}(\vec{X}_{i},...,\vec{X}_{i})\Big{)}+\vec{U}_{i},$ (10)
where $\vec{U}_{i}=\vec{0}$ for non controlled nodes and, for $i=i_{p}$, $\vec{U}_{i_{p}}=(u_{i_{p}}(t),u_{i_{p}}(t))^{\top}$, with
$u_{i_{p}}(t)=\lambda_{i_{p}}[\Theta(t)-\Theta(t-t_{p})],$ (11)
where $\lambda_{i_{p}}$ are parameters drawn from a uniform distribution of a given interval and $\Theta$ is the Heaviside step function, whose value is $1$ when the argument is positive and 00 when it is null or negative. This way, the control is active as long as $t\leq t_{p}$.
@@ II.3.2 Control protocol II: parametric pinning control
The second type of pinning control we are going to implement consists in acting on the controlled nodes by modifying the parameters of the dynamical system. This protocol, which we will call parametric pinning control, is given by the following equations
$\dot{\vec{X}}_{i}=\vec{F}_{i}(\vec{X}_{i})+\mathbb{D}\sum_{j_{1},...,j_{d}=1}^{N}A_{i,j_{1},...,j_{d}}^{(d)}\Big{(}\vec{H}(\vec{X}_{j_{1}},...,\vec{X}_{j_{d}})-\vec{H}^{(d)}(\vec{X}_{i},...,\vec{X}_{i})\Big{)},$ (12)
where $\vec{F}_{i}=\vec{F}$ for the nodes that are not pinned, while, for $i\equiv i_{p}=1,...,N_{p}$, it is given by $\vec{F}_{i}=(f_{i_{p}},g_{i_{p}})^{\top}$, where
$\begin{cases}f_{i_{p}}(x_{i_{p}},y_{i_{p}})=\alpha x_{i_{p}}(t)-\Omega_{i_{p}}(t)y_{i_{p}}(t)-\left(x_{i_{p}}^{2}(t)+y_{i_{p}}^{2}(t)\right)x_{i_{p}}(t),\\ \\ g_{i_{p}}(x_{i_{p}},y_{i_{p}})=\Omega_{i_{p}}(t)x_{i_{p}}(t)+\alpha y_{i_{p}}(t)-\left(x_{i_{p}}^{2}(t)+y_{i_{p}}^{2}(t)\right)y_{i_{p}}(t).\end{cases}$ (13)
The frequency of the controlled nodes, $\Omega_{i_{p}}(t)$, is given by
$\Omega_{i_{p}}(t)=\omega\Theta(t-t_{p})+\omega_{i_{p}}\Theta(t_{p}-t),$ (14)
where $\omega_{i_{p}}$ is new frequency induced by the pinning and $\Theta$ is the Heaviside step function, whose value is $1$ when the argument is positive and 00 when it is null or negative. This ensures that $\Omega_{i_{p}}=\omega_{i_{p}}$ for $t\leq t_{p}$ and it switches to $\omega$ as soon as the control is switched off. In our numerical implementation, the new frequencies $\omega_{i_{p}}$ will be drawn from a uniform distribution of a given positive interval. As a last remark, let us note that the control acts only on the frequency and not on the amplitude, because we have numerically verified that acting on the amplitude has no effect whatsoever.
@@ III Numerical Results
In this section we show the numerical results of our pinning approaches. We start by comparing the case of a $3$-hyperring, where chimera states occur with both pinning approaches, with its corresponding clique-projected network, where chimera states do not emerge. Then, we will exploit the hyperring structure and develop a pinning strategy that maximizes the width of the incoherence region while the number of controlled nodes remains low. For the latter, we will show the results only for additive pinning, leaving the discussion of the analogous results obtained through parametric pinning in Appendix B.
@@ III.1 Comparison between higher-order and pairwise interactions
Let us proceed in testing our pinning approaches to control the emergence of chimera states on a $3$-hyperring and compare it with the clique-project network case. In Figs. 3 and 4, we show the results for the additive pinning on a hyperring and a clique-projected network, respectively, while, in Figs. 5 and 6, we show the results for the parametric pinning on a hyperring and a clique-projected network, respectively. For both pinning approaches, we see that the control induces a chimera state when the topology is higher-order (Figs. 3 and 5). Moreover, such state persists for long integration times [^4] In the figures, we show the temporal evolution until $1000$ time steps, while the chimera persists until about $4000$ time steps. After such integration time, the chimera state turns into an incoherent state where variation between adjacent phases is smooth, but $\mathcal{R}_{i}^{\mathcal{E}}$ is low. We have found this state to persist with constant $\mathcal{R}_{i}^{\mathcal{E}}$ until $20000$ time steps (the maximum integration time tested).. These features can be visualized through the hyperedge-induced local order parameter $\mathcal{R}_{i}^{\mathcal{E}}$, which shows that the nodes which have been controlled are not oscillating coherently with respect to their neighbors sharing the same hyperedge(s) and that such incoherence persists. On the other hand, the pairwise case does not yield chimeras (Figs. 4 and 6). In fact, the initial decoherence induced by the control is quickly reabsorbed by the system and, although a clear trace of the pinning remains, the difference between the phases of adjacent oscillators is small and the variation smooth. This can be visualized through the clique-based order parameter $\mathcal{R}_{i}^{\mathcal{C}}$. Note that in Muolo et al. (2024) the latter state was distinguished from a chimera one through the total phase variation. Our approach based on a local order parameter is complementary.
@@
Figure 3: Additive pinning induces phase chimera states on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes. Panel a) depicts the whole time series of variables $y_{i}(t)$ with $i=1,...,N$, panel b) the hyperedge-based local order parameter, panel c) a snapshot of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of variables $y_{i}(t)$ with $i=1,...,N$. The model parameters are $\alpha=1$ and $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\lambda_{i_{p}}$ are drawn from a uniform distribution in the interval $[-2,2]$.
@@
Figure 4: Additive pinning on a clique-projected network of $204$ nodes. Chimera states do not emerge in this setting, at contrast with the previous Fig. 3. Panel a) depicts the whole time series of variables $y_{i}(t)$ with $i=1,...,N$, panel b) the clique-based local order parameter, panel c) a snapshot of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of variables $y_{i}(t)$ with $i=1,...,N$. The model parameters are $\alpha=1$ and $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\lambda_{i_{p}}$ are the same of the previous figure.
Additionally, let us stress that all results hereby shown, regardless of the order of the hyperring, coupling and pinning approach, are due to the higher-order topology and no chimera states are found when performing the simulations with the same setting but on the corresponding clique-projected network, exactly as in the figures shown in this section. Note, though, that there is one particular exception discussed in Appendix A, where the observed pattern is not due to the higher-order topology but due to the coupling configuration, and, in fact, it is found also in the corresponding pairwise system. Such results provide further evidence that higher-order interactions promote the presence of chimera states and are consistent with the existing literature Kundu and Ghosh (2022b); Li et al. (2023); Muolo et al. (2024).
@@
Figure 5: Parametric pinning induces phase chimera states on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes. Panel a) depicts the whole time series of variables $y_{i}(t)$ with $i=1,...,N$, panel b) the hyperedge-based local order parameter, panel c) a snapshot of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of variables $y_{i}(t)$ with $i=1,...,N$. The model parameters are $\alpha=1$ and $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\omega_{i_{p}}$ are drawn from a uniform distribution in the interval $[0.5,2.5]$.
@@
Figure 6: Parametric pinning on a clique-projected network of $204$ nodes. Chimera states do not emerge in this setting, at contrast with the previous Fig. 5. Panel a) depicts the whole time series of variables $y_{i}(t)$ with $i=1,...,N$, panel b) the clique-based local order parameter, panel c) a snapshot of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of variables $y_{i}(t)$ with $i=1,...,N$. The model parameters are $\alpha=1$ and $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\omega_{i_{p}}$ are the same of the previous figure.

### v2

@@ II.1 Stuart-Landau oscillators coupled via nonlocal hyperrings
$\begin{cases}\displaystyle\dot{x}_{i}=\alpha x_{i}-\omega y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)x_{i}+\varepsilon\sum_{j}^{N}A_{ij}\Big{(}x_{j}^{3}-x_{i}^{3}\Big{)},\\ \\ \displaystyle\dot{y}_{i}=\omega x_{i}+\alpha y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)y_{i}+\varepsilon\sum_{j}^{N}A_{ij}\Big{(}x_{j}^{3}-x_{i}^{3}\Big{)}.\end{cases}$ (7)
The equations for interactions of different orders can be constructed analogously by means of the coupling functions (5) (see also the SM of muolo2024phase ). Before proceeding any further, let us stress the intrinsic difference between pairwise and higher-order interactions by looking at the coupling functions given by the above equations. In the pairwise setting, the clique-projected network may seem to represent a group interaction; however, in Eqs. (7) it emerges that the units interact one-by-one, at striking difference with Eqs. (6), where they interact simultaneously in a group. Stated differently, the clique-projected network represents an ensemble of interactions in pairs, while the hyperring represents a different type of interaction that involves simultaneously all the nodes of the cliques and that cannot be decomposed as the sum of pairwise interactions [^4] One can visualize the difference between the two settings by thinking of a social science example. Let us consider a group of four people interacting via a messaging app: a clique would represent the case in which they are writing messages to each other separately, while a hyperedge represents the interactions via a group chat..
@@ II.2 Hyperedge-based local order parameter
To quantify the synchronization of an ensemble of oscillators it is common to use the Kuramoto order parameter kuramoto1975 , which gives a global measure of how much the oscillators are synchronized. However, chimera states involve the coexistence of coherent (i.e., synchronized) and incoherent (i.e., desynchronized) oscillators and the respective regions are localized. Hence, a global measure of the synchronization does not provide useful information on the chimera state. For this reason, scholars have proposed a local Kuramoto order parameter, which measures the synchronization in a given region of the network, by quantifying the differences between neighboring oscillators, as was done, for instance, in gambuzza2016pinning . In the case of hyperrings, a natural extension of the local order parameter would need to account for the synchronization inside each hyperedge rather than some arbitrary neighborhood. Partially inspired by a work by Shanahan shanahan2010metastable , where the order parameter is defined with respect to the communities of a network, we hereby define a hyperedge-based local order parameter as follows
@@
where $\mathrm{\imath}$ is the imaginary unit, $\vartheta_{j}$ is the (time evolving) phase of the $j$-th oscillator and $\mathcal{E}_{i}$ is the hyperedge(s) node $i$ is part of. By taking as example a $3$-hyperring, Fig. 1a), we can see that, if node $i$ is a junction node, then it is part of $2$ different hyperedges and will have $6$ neighboring nodes, while non-junction nodes will have only $3$ neighbors. From that, we can observe that non-junction nodes will have the same hyperedge-based local order parameter $\mathcal{R}_{i}^{\mathcal{E}}$. The number of nodes with the same $\mathcal{R}_{i}^{\mathcal{E}}$ in each hyperedge increases with the order of the hyperring: for instance, in $3$-hyperrings they will be $2$, while in $6$-hyperrings they will be $5$, and so on.
@@ II.3 Pinning control
Chimera states are often elusive patterns, emerging only for limited ranges of the parameters and specific initial conditions zakharova2020chimera . Higher-order interactions greatly enhance the possibility of observing such a behavior ghosh_chimera2 ; muolo2024phase ; ghosh_chimera_high-order ; zhang2021unified . Nonetheless, ad hoc initial conditions remain a fundamental prerequisite for the chimera to emerge. Our goal is, hence, to induce chimera states in settings where they would not spontaneously appear. For this, we put to use a popular technique in control theory, called pinning control, which consists in externally acting on a subset of the nodes to drive the dynamics of the whole ensemble of nodes towards a desired state grigoriev1997pinning ; sorrentino2007controllability , and has been successfully applied in the context of chimera states on networks gambuzza2016pinning . In our setting, the pinning will act as a perturbation on a subset of the nodes, with the goal of developing a region of incoherence, while leaving the unperturbed nodes in their synchronous state.
Figure 2: Scheme of the pinning control for a $3$-hyperring of $24$ nodes. The system starts in the synchronized state and the control input is applied to a subset $\mathcal{I}_{p}=\{1,\ldots,N_{p}\}$ of the nodes to induce the emergence of a chimera state.
In this work, we employ two distinct protocols to implement pinning control in the system, namely additive pinning and parametric pinning. The equations governing the controlled system under each protocol are presented in the following subsections. Here, we introduce two parameters that are common to both protocols, which serve to specify the nodes that are subject to control (i.e., pinned) and the time interval during which the control is active. Specifically, we indicate with $t_{p}$ the time during which the control is active, and with $N_{p}<N$ the number of pinned nodes. In addition, we denote the set of all nodes by =$\mathcal{I}=\{1,\ldots,N\}$, and indicate the set of pinned nodes by $\mathcal{I}_{p}$. In most cases, we assume that the pinned nodes correspond to the first $N_{p}$ nodes, i.e., $\mathcal{I}_{p}=\{1,\ldots,N_{p}\}$. The pinning control setting, under this assumption, is schematically depicted in Fig. 2.
@@ II.3.1 Control protocol I: additive pinning control
Let us now proceed with the first control protocol, namely additive pinning control. As a preliminary remark, we note that this control strategy was successfully employed in gambuzza2016pinning to induce chimera states in systems with pairwise interactions. However, achieving such states required controlling approximately half of the nodes. In Sec. III, where we present our numerical results, we highlight the significant advantage gained by applying this technique in systems with higher-order interactions.
To describe the additive pinning control, we first rewrite the system equations as follows
$\dot{\vec{X}}_{i}=\vec{F}(\vec{X}_{i})+\mathbb{D}\sum_{j_{1},...,j_{d}=1}^{N}A_{ij_{1}...j_{d}}^{(d)}\Big{(}\vec{H}(\vec{X}_{j_{1}},...,\vec{X}_{j_{d}})-\vec{H}^{(d)}(\vec{X}_{i},...,\vec{X}_{i})\Big{)}+\vec{U}_{i},$ (10)
where $\vec{U}_{i}$ is an additive term implementing the control action. This term is given by
${\vec{U}_{i}=\left\{\begin{array}[]{l}\vec{0},\quad i\in\mathcal{I}\setminus\mathcal{I}_{p}\\ (u_{i}(t),u_{i}(t))^{\top},\quad i\in\mathcal{I}_{p}\end{array}\right.}$ (11)
with
$u_{i}(t)=\lambda_{i}[\Theta(t)-\Theta(t-t_{p})],$ (12)
where $\lambda_{i}$ are parameters drawn from a uniform distribution of a given interval and $\Theta$ is the Heaviside step function, whose value is $1$ when the argument is positive and 00 when it is null or negative. This way, the control is active as long as $t\leq t_{p}$.
As an example, we consider the application of the protocol to the $3$-hyperring. In this case, the equations of the pinned nodes read
${\begin{cases}\displaystyle\dot{x}_{i}=\alpha x_{i}-\omega y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)x_{i}+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{ij_{1}j_{2}j_{3}}^{(3)}\Big{(}x_{j_{1}}x_{j_{2}}x_{j_{3}}-x_{i}^{3}\Big{)}+\lambda_{i}[\Theta(t)-\Theta(t-t_{p})],\\ \\ \displaystyle\dot{y}_{i}=\omega x_{i}+\alpha y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)y_{i}+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{ij_{1}j_{2}j_{3}}^{(3)}\Big{(}x_{j_{1}}x_{j_{2}}x_{j_{3}}-x_{i}^{3}\Big{)}+\lambda_{i}[\Theta(t)-\Theta(t-t_{p})],\end{cases}}$ (13)
with $i\in\mathcal{I}_{p}$.
The uncontrolled nodes ($i\in\mathcal{I}\setminus\mathcal{I}_{p}$) are, instead, described by Eqs. (6).
@@ II.3.2 Control protocol II: parametric pinning control
In the second type of control, which we will call parametric pinning control, the control law acts on the pinned nodes by modifying the parameters of the dynamical system associated to each node. To formalize this protocol, we rewrite the system of equations as follows
$\dot{\vec{X}}_{i}=\vec{F}_{i}(\vec{X}_{i})+\mathbb{D}\sum_{j_{1},...,j_{d}=1}^{N}A_{ij_{1}...j_{d}}^{(d)}\Big{(}\vec{H}(\vec{X}_{j_{1}},...,\vec{X}_{j_{d}})-\vec{H}^{(d)}(\vec{X}_{i},...,\vec{X}_{i})\Big{)},$ (14)
where the uncoupled (or natural) dynamics of each node, namely $\vec{F}_{i}$, differentiates between pinned and unpinned nodes. This term is, in fact, given by
${\vec{F}_{i}=\left\{\begin{array}[]{l}\vec{F},\quad i\in\mathcal{I}\setminus\mathcal{I}_{p}\\ (f_{p},g_{p})^{\top},\quad i\in\mathcal{I}_{p}\end{array}\right.}$ (15)
where
${\begin{cases}f_{p}(x_{i},y_{i})=\alpha x_{i}(t)-\Omega_{p,i}(t)y_{i}(t)-\left(x_{i}^{2}(t)+y_{i}^{2}(t)\right)x_{i}(t),\\ \\ g_{p}(x_{i},y_{i})=\Omega_{p,i}(t)x_{i}(t)+\alpha y_{i}(t)-\left(x_{i}^{2}(t)+y_{i}^{2}(t)\right)y_{i}(t).\end{cases}}$ (16)
Here, the frequencies of the controlled nodes, $\Omega_{p,i}(t)$ for $i\in\mathcal{I}_{p}$, are given by
${\Omega_{p,i}(t)=\left\{\begin{array}[]{l}\omega_{p,i},\quad t\leq t_{p},\\ \omega,\quad t>t_{p},\end{array}\right.}$ (17)
where $\omega_{p,i}$ are new frequencies induced by the pinning; namely, we have that during the period where the control is on ($t\leq t_{p}$), $\Omega_{p,i}=\omega_{p,i}$, while $\Omega_{p,i}$ switches to $\omega$ as soon as the control is switched off. In our numerical implementation, the new frequencies $\omega_{p,i}$ will be drawn from a uniform distribution of a given positive interval. Note that, despite the pinning being a perturbation of the synchronous solution, in our simulations the order of magnitude of such perturbation is large, hence a linear stability analysis would not give any useful information. In fact, the latter being a local analysis, it requires small perturbations about the equilibrium state and it is not informative when the perturbations are large. Let us also note that the control acts only on the frequency and not on the amplitude, because we have numerically verified that acting on the amplitude has no effect whatsoever. This can be very well understood by looking at the phase reduction analysis of Sec. 4.
Lastly, we exemplify the parametric pinning control protocol by considering again a $3$-hyperring. The equations of the pinned nodes read
${\begin{cases}\displaystyle\dot{x}_{i}=\alpha x_{i}-\Omega_{p,i}y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)x_{i}+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{ij_{1}j_{2}j_{3}}^{(3)}\Big{(}x_{j_{1}}x_{j_{2}}x_{j_{3}}-x_{i}^{3}\Big{)},\\ \\ \displaystyle\dot{y}_{i}=\Omega_{p,i}x_{i}+\alpha y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)y_{i}+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{ij_{1}j_{2}j_{3}}^{(3)}\Big{(}x_{j_{1}}x_{j_{2}}x_{j_{3}}-x_{i}^{3}\Big{)},\end{cases}}$ (18)
with $i\in\mathcal{I}_{p}$ and $\Omega_{p,i}$ as in Eq. (17).
Also in this case, the uncontrolled nodes ($i\in\mathcal{I}\setminus\mathcal{I}_{p}$) are, instead, described by Eqs. (6).
@@ III Numerical results
In this section we show the numerical results of our pinning approaches. We start by considering the case $\mathcal{I}_{p}=\{1,\ldots,N_{p}\}$, and compare the results obtained on a $3$-hyperring, where chimera states occur with both pinning approaches, with its corresponding clique-projected network, where chimera states do not emerge. Then, by leveraging the structure of the hyperring, we develop a different strategy to select the pinned nodes, i.e., we define a different set $\mathcal{I}_{p}$ based on the structure under control, and show how this maximizes the width of the incoherence region while the number of controlled nodes remains low. For the latter, we will show the results only for additive pinning, leaving the discussion of the analogous results obtained through parametric pinning in Appendix B.
Lastly, let us point out that there is no universal consensus on when a state is chimera or not. There are some obvious features and characteristics, both quantitative and phenomenological, however, scholars may give different definition to patterns exhibiting some incoherence haugland2021changing . Here, we will call chimeras the states where the incoherence regions exhibit the chaotic behavior shown by Kuramoto and Battogtokh kuramoto2002coexistence and, for our specific case of phase chimeras, by Zajdela and Abrams zajdela_abrams . We will call states exhibiting a weak incoherence weak chimera-like states, inspired by the definition given by Aswhin and Burylko for chimera states with respect to frequencies ashwin2015weak .
@@ III.1 Comparison between higher-order and pairwise interactions
Figure 3: Additive pinning induces phase chimera states on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes. Panel a) depicts the whole time series of variables $y_{i}(t)$ with $i=1,...,N$, panel b) the hyperedge-based local order parameter, panel c) a snapshot of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of variables $y_{i}(t)$ with $i=1,...,N$. The results for the variables $x_{i}(t)$ are analogous and, hence, not shown. The model parameters are $\alpha=1$ and $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\lambda_{i_{p}}$ are drawn from a uniform distribution in the interval $[-2,2]$.
@@
Figure 4: Additive pinning on a clique-projected network of $204$ nodes. Chimera states do not emerge in this setting, at contrast with the previous Fig. 3. Panel a) depicts the whole time series of variables $y_{i}(t)$ with $i=1,...,N$, panel b) the clique-based local order parameter, panel c) a snapshot of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of variables $y_{i}(t)$ with $i=1,...,N$. The results for the variables $x_{i}(t)$ are analogous and, hence, not shown. The model parameters are $\alpha=1$ and $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\lambda_{i_{p}}$ are the same of the previous figure.
@@
Figure 5: Parametric pinning induces phase chimera states on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes. Panel a) depicts the whole time series of variables $y_{i}(t)$ with $i=1,...,N$, panel b) the hyperedge-based local order parameter, panel c) a snapshot of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of variables $y_{i}(t)$ with $i=1,...,N$. The results for the variables $x_{i}(t)$ are analogous and, hence, not shown. The model parameters are $\alpha=1$ and $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\omega_{i_{p}}$ are drawn from a uniform distribution in the interval $[0.5,2.5]$.
@@
Figure 6: Parametric pinning on a clique-projected network of $204$ nodes. Chimera states do not emerge in this setting, at contrast with the previous Fig. 5. Panel a) depicts the whole time series of variables $y_{i}(t)$ with $i=1,...,N$, panel b) the clique-based local order parameter, panel c) a snapshot of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of variables $y_{i}(t)$ with $i=1,...,N$. The results for the variables $x_{i}(t)$ are analogous and, hence, not shown. The model parameters are $\alpha=1$ and $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\omega_{i_{p}}$ are the same of the previous figure.
Let us proceed in testing our pinning approaches to control the emergence of chimera states on a $3$-hyperring and compare it with the clique-project network case. In Figs. 3 and 4, we show the results for the additive pinning on a hyperring and a clique-projected network, respectively, while, in Figs. 5 and 6, we show the results for the parametric pinning on a hyperring and a clique-projected network, respectively. Here we always assume $\mathcal{I}_{p}=\{1,\ldots,N_{p}\}$. Note that the time series is depicted for the $y$ variables of the SL model, but the results are qualitatively the same if we display the $x$ variable; indeed, the local order parameter is computed using both variables. For both pinning approaches, we see that the control induces a chimera state when the topology is higher-order (Figs. 3 and 5). This can be visualized through the hyperedge-based local order parameter $\mathcal{R}_{i}^{\mathcal{E}}$, which shows that the nodes that have been controlled are not oscillating coherently with respect to their neighbors sharing the same hyperedge(s) and that such incoherence persists. On the other hand, the pairwise case does not yield chimeras (Figs. 4 and 6). In fact, the initial decoherence induced by the control is quickly reabsorbed by the system and, although a clear trace of the pinning remains, the difference between the phases of adjacent oscillators is small and the variation smooth. This can be visualized through the clique-based order parameter $\mathcal{R}_{i}^{\mathcal{C}}$. Let us point out that clique-projected networks obtained from hyperrings of different orders do not significantly affect the behavior of the system. Here, we have shown the results on a clique-projected network as depicted in Fig. 1, but qualitatively similar time series are obtained on different clique-projected networks, which is consistent with the results obtained in a previous work muolo2024phase . Moreover, when the nonlinear coupling have the same functional form, the results on different clique-projected networks are the same. Note that in muolo2024phase the dynamics on clique-projected networks was distinguished from a chimera one through the total phase variation. Our approach based on a local order parameter is complementary.
For the chosen parameters, the higher-order setting always yields chimera states when the consecutive pinned nodes are about $10$ and above. For lower numbers of pinned nodes, we will always have at least a weak chimera-like state, while the emergence of chimera states depends on the random perturbation of the pinning. Note that, when the number of pinned nodes is very small (e.g., $1\sim 3$), the resulting pattern will always be a weak chimera-like state. Moreover, such weak chimera-like state persists for long integration times. In Figs. 3 and 5, we show the temporal evolution until $1000$ time units, while the chimera state persists until about $3000\sim 4000$ time units (depending on the perturbation of the pinning). After such integration time, the chimera state turns into a weak chimera-like state, where variation between adjacent phases is smoother, but $\mathcal{R}_{i}^{\mathcal{E}}$ remains low. We have found that the weak chimera state persists with almost constant $\mathcal{R}_{i}^{\mathcal{E}}$ until $20000$ time units (the maximum integration time tested). This is particularly striking when compared to the behavior observed on clique-projected networks, where the incoherence vanishes after a short time. To better appreciate the difference between the higher-order and pairwise cases, we provide $4$ Supplementary Videos (SVs), where the dynamics until $5000$ time units can be visualized: SV1 is obtained from a setting analogous to Fig. 3 (i.e., additive pinning on a $3$-hyperring), SV2 is obtained from a setting analogous to Fig. 4 (i.e., additive pinning on a clique-projected network), SV3 is obtained from a setting analogous to Fig. 5 (i.e., parametric pinning on a $3$-hyperring), and SV4 is obtained from a setting analogous to Fig. 6 (i.e., parametric pinning on a clique-projected network).
Additionally, let us stress that all results hereby shown, regardless of the order of the hyperring, coupling and pinning approach, are due to the higher-order topology and no chimera states are found when performing the simulations with the same setting but on the corresponding clique-projected network, exactly as in the figures shown in this section. Note, though, that there is one particular exception discussed in Appendix A, where the observed pattern is not due to the higher-order topology but due to the coupling configuration, and, in fact, it is found also in the corresponding pairwise system. Such results provide further evidence that higher-order interactions promote the presence of chimera states and are consistent with the existing literature ghosh_chimera2 ; muolo2024phase ; ghosh_chimera_high-order ; zhang2021unified .